# RepoHyper: Better Context Retrieval Is All You Need for Repository-Level Code Completion

### v2

@@ Abstract
Code Large Language Models (CodeLLMs) have demonstrated impressive proficiency in code completion tasks. However, they often fall short of fully understanding the extensive context of a project repository, such as the intricacies of relevant files and class hierarchies, which can result in less precise completions. To overcome these limitations, we present RepoHyper, a multifaceted framework designed to address the complex challenges associated with repository-level code completion. Central to RepoHyper is the Repo-level Semantic Graph (RSG), a novel semantic graph structure that encapsulates the vast context of code repositories. Furthermore, RepoHyper leverages Expand and Refine retrieval method, including a graph expansion and a link prediction algorithm applied to the RSG, enabling the effective retrieval and prioritization of relevant code snippets. Our evaluations show that RepoHyper markedly outperforms existing techniques in repository-level code completion, showcasing enhanced accuracy across various datasets when compared to several strong baselines. Our implementation of RepoHyper can be found at https://github.com/FSoft-AI4Code/RepoHyper.
RepoHyper: Better Context Retrieval Is All You Need
for Repository-Level Code Completion
Huy N. ${}^{\spadesuit}$  Hoang N. ${}^{\clubsuit}$  Tien N. ${}^{\diamondsuit}$ Nghi D. Q. ${}^{\heartsuit}$$^{\spadesuit}$ Software AI Center, Viet Nam ${}^{\clubsuit}$epartment of Computer Science, Nanyang Technological University, Singapore ${}^{\diamondsuit}$omputer Science Department, The University of Texas at Dallas, USA ${}^{\heartsuit}$ulbright University, Viet Nam huypn16@fpt.com, C210055@ntu.edu.sg, tien.n.nguyen@utdallas.edu, nghi.bui@fulbright.edu.vn
@@ 2.1 Code LLMs for code generation & understanding
Recent research has introduced a plethora of Large Language Models (LLMs) tailored for code-related tasks Bui et al. (2023), aiming to enhance code understanding and generation. These models are categorized into closed-source and open-source variants. Initially, closed-source models like Codex Chen et al. (2021a), Code-Davinci Chen et al. (2021a), and PaLM-Coder Chowdhery et al. (2023) demonstrated exceptional performance on well-known code completion benchmarks, including HumanEval Chen et al. (2021a), MBPP Austin et al. (2021), and APPS Hendrycks et al. (2021). Subsequently, the emergence of open-source models such as the CodeGen series Nijkamp et al. (2023b, a), CodeT5 Wang et al. (2021a), CodeT5+ Wang et al. (2023), CodeGeeX Zheng et al. (2023), StarCoder Li et al. (2023), Wizard Coder Luo et al. (2023), CodeLlama Roziere et al. (2023), and DeepSeek-Coder Guo et al. (2024) began to rival the closed-source models in terms of benchmark performance. Despite their purported efficacy across a broad spectrum of code intelligence tasks, code generation and completion emerge as their most notable and widely utilized applications. Nonetheless, a common limitation among these models is their inability to grasp the extensive contexts of entire code repositories, constrained by token count limits. Moreover, their training predominantly on public GitHub codebases overlooks the nuanced local context specific to individual repositories. Consequently, their performance tends to falter in repository-level completion tasks, as highlighted in recent evaluations Liu et al. (2023a).
@@ 2.2 Repository-level Code Completion
Research in repository-level code completion has seen notable advancements through works like RLPG Wang et al. (2021b), CoCoMIC Ding et al. (2022), RepoCoder Liu et al. (2023b), CodePlan Bairi et al. (2023) and A3-Codegen Liao et al. (2023). These studies highlight the critical role of leveraging both within-file and cross-file contexts to improve code completion accuracy. Conversely, initiatives such as RepoFusion Shrivastava et al. (2023a), RepoPrompts Shrivastava et al. (2023b), and MGD Agrawal et al. (2023) delve into methodologies for effectively integrating these contexts, assuming their availability from external sources. Additionally, contributions like RepoBench Liu et al. (2023a) and CrossCodeEval Ding et al. (2023) emphasize the need for specialized benchmarks designed to evaluate code completion systems within the broader, repository-level contexts, aiming for a comprehensive capability assessment. Furthermore, CodeAgent Zhang et al. (2024) introduces an innovative benchmark incorporating documentation, contexts, runtime environments, and a pipeline for interacting with repositories through multi-agent systems.
@@ 3 Methodology
Figure 2 illustrates the overall architecture of our approach, RepoHyper. Given an existing, incomplete code snippet $Q$, we first encode it into a semantic vector using an encoder function:
$Z_{Q}=\text{Encoder}(Q).$
Our objective is to retrieve relevant semantic contexts $T$ from the repository $R$, which is formulated as:
$\text{T}=\text{Retriever}(Q,\text{R}).$
These retrieved contexts are subsequently utilized by a Large Language Model (LLM) to generate the final code prediction:
@@ 3.2 Search-then-Expand Strategies
where PATH denotes the type of BFS exploration path from kNN searched node $V_{j}\in A$ to any walking node $V_{i}$. For example, one might prioritize the exploration from a class node to its script node then to imported function to draw a possible invoke relationship for code completion (class-script-function path type), but one is unlikely to explore chain of method calls (method-method-method path type).
@@ 3.3 Re-ranking as Link Prediction
In all experiments, we employ GraphSAGE Hamilton et al. (2017) with $L$ layers as the graph neural network (GNN) model to update the representation for each node based on the passage graph. The $l$-th layer of the GNN model updates the embedding of node $i$ as follows:
@@ 4 Empirical Evaluation
Table 1: Results of RepoHyper on RepoBench-R dataset. The studied encoder models include codebert-base for CodeBERT, unixcoder-base for UniXcoder, codet5p for using CodeT5+ 770M parameters. RepoHyper has UniXcoder and CodeT5+ 770M as the base encoders, respectively. Numbers are shown in percentage (%), with the best performance highlighted in bold.
@@ 4.2.1 Context Retrieval on RepoBench-R
We follow  Liu et al. (2023a) to define 4 baselines: (1) Random Retrieval, where code snippets are chosen randomly, providing a basic comparison level. This process is repeated 100 times to average the results for consistency. (2) Lexical Retrieval, which uses simple text comparison techniques like Jaccard Similarity and Edit Distance to find relevant snippets based on the code tokens. (3) Similarity-based Retrieval, employing encoder models like CodeBERT and UnixCoder to generate code embeddings and Cosine Similarity to measure the semantic similarity between the cropped code and the candidate snippets. In RepoHyper setting, we employ our pipeline, which leverages different encoder models (UniXCoder Guo et al. (2022), CodeT5+-770M Wang et al. (2023)) to encode the query, construct semantic graph, node expansion, and link prediction to retrieve the contexts.
@@ 5.2 Performance on Code Completion
Gpt-3.5-turbo ${}^{\ast}$ 26.35 33.14 36.31 44.01
${}^{\ast}$ 30.59 38.37 40.65 49.12
@@
DeepSeek-Coder ${}^{\ast}$ 25.46 32.39 34.23 42.13
${}^{\ast}$ 27.29 35.15 37.10 46.08
@@ 6.2 Retrieved Nodes Analysis
ptSimNptParClsptChiClsptSibptImpSibptImpSimNptImpParCls000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Context TypesRetrieval Performance Acc@5RepoHyperSimilarity-based Search
@@ 6.3 Ablation Study
We vary different components of our model to understand their contributions, including the Search-then-Expand component and the Link Predictor. As a case study, we take the code retrieval task for experiment and present the results in Table 4.
Several meaningful observations can be drawn: Without Link Predictor, Pattern Search offers better accuracy than Exhausted Search. This is likely due to the fact that Exhausted Search has to include more nodes to obtain similar hits rate making the later re-ranking more difficult because of the noisy context nodes. Using Re-ranking loss shows inferior performance compared with Link Predictor, we did some analysis and found that using Re-ranking approach causes over-smoothing, thus damages decision boundary for the most optimal context.
@@ 7 Conclusion
In this paper, we introduced RepoHyper, a novel framework aimed at enhancing repository-level code completion by addressing its complex challenges. RepoHyper advances this domain through three key components: the Repo-level Semantic Graph (RSG), Search-then-Expand Strategies, and a Link Predictor, collectively improving the accuracy and relevance of code suggestions.
Our evaluations, conducted on the RepoBench benchmark, demonstrate RepoHyper’s superior performance over state-of-the-art methods.
@@ A.1 Building RSG
A.2 Context Types in Retrieved Node Analysis
@@
A.3 RepoBench Benchmark Statistics
@@
A.4 Samples

### v4

# RepoHYPER: : Search-Expand-Refine on Semantic Graphs for Repository-Level Code Completion

@@ Abstract
Code Large Language Models (CodeLLMs) have demonstrated impressive proficiency in code completion tasks. However, they often fall short of fully understanding the extensive context of a project repository, such as the intricacies of relevant files and class hierarchies, which can result in less precise completions. To overcome these limitations, we present RepoHyper, a multifaceted framework designed to address the complex challenges associated with repository-level code completion. Central to RepoHyper is the Repo-level Semantic Graph (RSG), a novel semantic graph structure that encapsulates the vast context of code repositories. Furthermore, RepoHyper leverages Expand and Refine retrieval method, including a graph expansion and a link prediction algorithm applied to the RSG, enabling the effective retrieval and prioritization of relevant code snippets. Our evaluations show that RepoHyper markedly outperforms existing techniques in repository-level code completion, showcasing enhanced accuracy across various datasets when compared to several strong baselines. Our implementation is published at.  [^1] https://anonymous.4open.science/r/RepoHyper-3836/README.md
RepoHYPER: : Search-Expand-Refine on Semantic Graphs for
Repository-Level Code Completion
Huy N. Phan  Hoang N. Phan  Tien N. Nguyen Nghi D. Q. Bui FPT Software AI Center, Viet Nam Department of Computer Science, Nanyang Technological University, Singapore Computer Science Department, The University of Texas at Dallas, USA Fulbright University, Viet Nam huypn16@fpt.com, C210055@ntu.edu.sg, tien.n.nguyen@utdallas.edu, nghi.bui@fulbright.edu.vn
@@ 2.1 Code LLMs for code generation & understanding
Recent research has introduced a plethora of Large Language Models (LLMs) tailored for code-related tasks Bui et al. (2023); Chowdhery et al. (2023); Chen et al. (2021a); Austin et al. (2021); Hendrycks et al. (2021); Nijkamp et al. (2023b, a); Zheng et al. (2023); Wang et al. (2023); Bui and Jiang (2018); Jayasundara et al. (2019); Bui et al. (2023); Guo et al. (2024); Li et al. (2023); Roziere et al. (2023), aiming to enhance code understanding and generation. These models are categorized into closed-source and open-source variants. Initially, closed-source models like Codex Chen et al. (2021a), Code-Davinci Chen et al. (2021a), and PaLM-Coder Chowdhery et al. (2023) demonstrated exceptional performance on well-known code completion benchmarks, including HumanEval Chen et al. (2021a), MBPP Austin et al. (2021), and APPS Hendrycks et al. (2021). Subsequently, the emergence of open-source models such as the CodeGen series Nijkamp et al. (2023b, a), CodeT5 Wang et al. (2021a), CodeT5+ Wang et al. (2023), CodeGeeX Zheng et al. (2023), StarCoder Li et al. (2023), Wizard Coder Luo et al. (2023), CodeLlama Roziere et al. (2023), and DeepSeek-Coder Guo et al. (2024) began to rival the closed-source models in terms of benchmark performance. Despite their purported efficacy across a broad spectrum of code intelligence tasks, code generation and completion emerge as their most notable and widely utilized applications.
@@ 2.2 Repository-level Code Completion
Repository-level code completion has seen notable advancements through works like RLPG Wang et al. (2021b), CoCoMIC Ding et al. (2022), RepoCoder Liu et al. (2023b), CodePlan Bairi et al. (2023) and A3-Codegen Liao et al. (2023). These studies highlight the critical role of leveraging both within-file and cross-file contexts to improve code completion accuracy. Conversely, RepoFusion Shrivastava et al. (2023a), RepoPrompts Shrivastava et al. (2023b), and MGD Agrawal et al. (2023) propose methodologies for effectively integrating these contexts, assuming their availability from external sources. RepoBench Liu et al. (2023a) and CrossCodeEval Ding et al. (2023) emphasize the need for end-to-end benchmarks designed to evaluate code completion systems within the broader, repository-level contexts. Furthermore, CodeAgent Zhang et al. (2024) incorporates documentation, contexts, runtime environments, and a pipeline for interacting with repositories through multi-agent systems.
@@ 3 Methodology
Figure 2 illustrates the overall architecture of our approach, RepoHyper. Given an existing, incomplete code snippet $Q$, we first encode it into a semantic vector using an encoder function. Our objective is to retrieve relevant semantic contexts $T$ from the repository $R$. Consequently, these retrieved contexts are subsequently utilized by a Large Language Model (LLM) to generate the final code prediction:
@@ 3.2 Search-then-Expand Strategies
where PATH denotes the type of BFS exploration path from kNN searched node $V_{j}\in A$ to any walking node $V_{i}$. For example, one might prioritize the exploration from a class node to its script node then to imported function to draw a possible invoke relationship for code completion (class-script-function path type), but one is unlikely to explore chain of method calls (method-method-method path type). More details are in Appendix A.3.
@@ 3.3 Re-ranking as Link Prediction
More details about incorporating query node into inference process can be found in Appendix. A.6. In all experiments, we employ GraphSAGE Hamilton et al. (2017) with $L$ layers as the graph neural network (GNN) model to update the representation for each node based on the passage graph. The $l$-th layer of the GNN model updates the embedding of node $i$ as follows:
@@ 4 Empirical Evaluation
Table 1: Results of RepoHyper on RepoBench-R dataset. The studied encoder models include codebert-base for CodeBERT, unixcoder-base for UniXcoder, codet5p for using CodeT5+ 770M parameters. RepoHyper has UniXcoder and CodeT5+ 770M as the base encoders, respectively. Numbers are shown in percentage (%), with the best performance highlighted in bold. text-embedding-large-3 for OpenAI.
@@ 4.2.1 Context Retrieval on RepoBench-R
We follow  Liu et al. (2023a) to define 4 baselines: (1) Random Retrieval, where code snippets are chosen randomly, providing a basic comparison level. This process is repeated 100 times to average the results for consistency. (2) Lexical Retrieval, which uses simple text comparison techniques like Jaccard Similarity and Edit Distance to find relevant snippets based on the code tokens. (3) Similarity-based Retrieval, employing encoder models like CodeBERT, UnixCoder and OpenAI Text Embedding Models to generate code embeddings and Cosine Similarity to measure the semantic similarity between the cropped code and the candidate snippets. In RepoHyper setting, we employ our pipeline, which leverages different encoder models (UniXCoder Guo et al. (2022), CodeT5+-770M Wang et al. (2023)) to encode the query, construct semantic graph, node expansion, and link prediction to retrieve the contexts.
@@ 5.2 Performance on Code Completion
Gpt-3.5-turbo In-File-Only 26.35 33.14 36.31 44.01
Gold-Only 30.59 38.37 40.65 49.12
@@
DeepSeek-Coder In-File-Only 25.46 32.39 34.23 42.13
Gold-Only 27.29 35.15 37.10 46.08
@@ 6.2 Retrieved Nodes Analysis
SimNParClsChiClsSibImpSibImpSimNImpParCls000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Context TypesRetrieval Performance Acc@5RepoHyperSimilarity-based Search
@@ 6.3 Ablation Study
Several meaningful observations can be drawn from Table 4: without Link Predictor, Pattern Search offers better accuracy than Exhausted Search. This is likely due to the fact that Exhausted Search has to include more nodes to obtain similar hits rate making the later re-ranking more difficult because of the noisy context nodes.
@@ 7 Conclusion
In this paper, we introduced RepoHyper, a novel framework aimed at enhancing repository-level code completion by addressing its complex challenges. RepoHyper advances this domain through three key components: the Repo-level Semantic Graph (RSG), Search-then-Expand Strategies, and a Link Predictor, collectively improving the accuracy and relevance of code suggestions. With extensive evaluation, RepoHyper show superiority in Repo-level Code Completion.
@@ A.1 Building RSG
A.2 Gold Snippet Definition
In the training set of RepoBench-R, for each sample, every snippet parsed from import statements is treated as a potential candidate for next-line prediction, for example, we have definitions of two following imported functions: add and minus as the snippets and the gold snippet is the optimal context for prediction, here which is "add" from src import add, minus nextline: add(1,2).
A.3 Pattern Search
In Repository Semantic Graph (RSG), there are three types of nodes: function (method), class, and script. The edges between these nodes represent five different relationships: import (1), invoke (2), ownership (3), encapsulate (4), and hierarchy (5). A path type is defined by the sequence of edge types along that path in the RSG. For example, let’s consider a node $V_{j}$ identified by the $k$-nearest neighbor (kNN) search, and a target node $V_{\text{target}}$ representing the “gold snippet" (e.g., a method called by $V_{j}$). There can be multiple paths from $V_{j}$ to $V_{\text{target}}$. One such path could be: $V_{j}$ (method) $\rightarrow$ (owned by) $V_{1}$ (class) $\rightarrow$ (encapsulated in) $V_{2}$ (script) $\rightarrow$ (import) $V_{\text{target}}$ This path has a type of (ownership, encapsulate, import), represented by the edge types (3, 4, 1). However, since there are multiple paths from $V_{j}$ to $V_{\text{target}}$, including all the nodes produced by these paths might introduce irrelevant contexts into the subgraph. For example, $V_{j}$ can reach $V_{\text{target}}$ through a more useful path like: $V_{j}$ (method) $\rightarrow$ (owned by) $V_{1}$rightarrow (owns) $V_{3}$ (method) $\rightarrow$ (call) $V_{\text{target}}$ This path is more useful because $V_{3}$ (method) can provide hints on how to call $V_{\text{target}}$, whereas the previous path included $V_{2}$ (script), which is a longer context and less useful for understanding how to use $V_{\text{target}}$. In the training set of RepoBench-R, for each sample, we have a gold snippet and an in-file context. We first use kNN search to find the $k$ most similar context nodes to the in-file context. Then, we perform an exhaustive search from these identified nodes to collect all possible paths leading to the gold snippet. For example, in sample 1, we might collect paths of types (3, 4, 1) and (1, 2, 3), while in sample 2, we might find paths of types (3, 4, 1) and (2, 4). The most frequent path type in the training set is (3, 4, 1), which appears twice in this example.
A.4 Context Types in Retrieved Node Analysis
@@
A.5 RepoBench Benchmark Statistics
@@
A.6 Rerank as Link Prediction
After extracting the subgraph using the k-nearest neighbor (kNN) search and expansion strategy, it is crucial to incorporate the query node, representing the in-file context, into the subgraph. To enable meaningful message passing between the query node and the extracted subgraph nodes, we need to establish edges connecting the query node to the subgraph. Fortuitously, in most cases, the query node (in-file context) already possesses certain connections to other nodes within the subgraph. For instance, the in-file context might invoke some functions that are represented as nodes in the subgraph for the task of next line prediction. By including the query node and its existing connections to the subgraph, the message passing network can effectively leverage the in-file context information during the link prediction process, consequently leading to more accurate and relevant code completion suggestions. This integration of the query node into the subgraph allows for a comprehensive understanding of the code context, facilitating the message passing network to make informed predictions. By considering both the subgraph extracted from the repository and the in-file context, the model can capture the relevant relationships and dependencies, resulting in code completion suggestions that are tailored to the specific code context at hand.
A.7 Samples